OVHCloud Network Status

Current status
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#13840 — BHS <> NWK
Scheduled Maintenance Report for Network & Infrastructure
We're continuing to put the 100G links into production.

The 20*10G links between the Beauharnois datacentre (Quebec, Canada) and the POP in Newark (New Jersey, USA) will be replaced by 6*100G.


Date: 2015-07-23 09:00:21 UTC
The flapping has now stopped. We will remount these links.

Date: 2015-07-02 07:55:09 UTC
There are 2 breaks in the links.

Work is in progress. We will postpone the job until everything is stable.

Date: 2015-07-02 07:54:18 UTC
Done. Monitoring the traffic.

Date: 2015-07-01 15:02:41 UTC
We're still migrating the traffic over.
Posted Jul 01, 2015 - 14:45 UTC